# AeroVerse: UAV-Agent Benchmark Suite for Simulating, Pre-training, Finetuning, and Evaluating Aerospace Embodied World Models

### v1

@@ V Dataset Suite
Figure 4: The left and right panels illustrate the construction schemes and statistics of the AerialAgent-Ego10k and CyberAgent-Ego500k datasets, respectively.
@@
V-A AerialAgent-Ego10k
Multi-Resolution UAV First-Person View City Images. The first-person view images of real cities captured by drones are derived from the UrbanBIS dataset, which is collected using aerial photogrammetry and encompasses a wide array of urban scenes. Specifically, the UrbanBIS dataset[31] comprises $0.5$ TB of aerial photographs from six actual locations: Qingdao, Wuhu, Longhua, Yuehai, Lihu, and Yingrenshi, covering a significant urban area of $10.78$ km² and including $3,370$ buildings, with a total of $113,346$ aerial photogrammetry images. We have requested images from the authors for the regions of Lihu, Longhua, Yingrenshi, and Yuehai, with resolutions of $6000\times 4000$, $8192\times 5640$, $5472\times 3648$, and $5472\times 3648$, respectively, yielding a total of $15,094$ images. From this dataset, we randomly selected $10,000$ images to serve as first-person view representations of real cities captured by drones.
Fine-grained Multi-attribute First-view Text Generation. To generate high-quality environmental descriptions, we utilize LLaVA-1.5-13B[32] to produce detailed accounts of surrounding buildings, roads, trees, and other scenery from first-person perspective images captured by a drone, as illustrated in Figure 4 left (a). To standardize the format of the environmental descriptions generated by LLaVA-1.5-13B[32], we employ specific prompts that emphasize the quantity, appearance, and shape of the buildings in the images, particularly focusing on the spatial relationships among the objects. This approach enhances the spatial reasoning capabilities of the drone agent. Furthermore, we specify that the sky should not be described, as this scene is relatively uniform and appears consistent from various perspectives of the drone, providing insufficient information. Consequently, the generated descriptions ensure a degree of diversity, accuracy, and detail.
Diverse Data Distribution. We perform a quantitative statistical analysis on AerialAgent-Ego10k. Figure 4 (b) and (c) illustrate the probability density functions (PDFs) of text vocabulary length and text sentence length, respectively, both exhibiting a shape akin to a normal distribution. This finding supports the rationality of the text distribution. The maximum length for image descriptions is $440$ words, with an average length of $144$ words. The maximum number of sentences in image descriptions is $42$, with an average of $11$ sentences per image. Both the number of sentences and text lengths exceed those of most existing visual-language datasets. Figure 4 (d) reveals that the dataset contains a total of $158,379$ sentences and $2,167,455$ words, of which $21,489$ are unique.
@@ V-B CyberAgent-Ego500k
First-Person Image-Text-Pose Generation. As illustrated on the right side of Figure 4 (a), the dataset construction method aligns with that of AerialAgent-10k and exhibits the following three characteristics:
@@ V-C SkyAgent-Scene3k
Dataset statistics. Figure 5 (b), (c), and (d) illustrate the distribution of description lengths, the number of sentences, and statistical information regarding scene descriptions. As shown in Figure 5 (b), the lengths of the descriptions range from $30$ to $80$ words. Generally, longer descriptions suggest a more complex scene with a greater number of environmental elements that require articulation. Figure 5 (c) indicates that most descriptions consist of four sentences, as we instruct annotators to depict each scene from four perspectives: front, back, left, and right. Descriptions containing 1$\sim$3 sentences occur when annotators consolidate multiple perspectives into a single sentence. In total, this dataset comprises $121,252$ words and $1,162$ distinct word types.
@@ V-D SkyAgent-Reason3k
Figure 6: The SkyAgent-Reason3k dataset contains $6$ types of reasoning patterns and corresponding concrete examples, as well as some statistical results.
@@ V-G SkyAgent-Act3k
Figure 9: Examples of drone agent actions in $2$ city scenarios (i.e., ShangHai and School) in SkyAgent-Act3k dataset, along with some instructions and statistics.
@@ VI-A Baselines
Baselines Selection. Due to the current scarcity of research on aerospace-embodied world models, we evaluate several mainstream and representative 3D and 2D visual-language models. This assessment aims to explore their potential and limitations concerning the proposed aerospace-embodied downstream task datasets, thereby providing a preliminary foundation for future researchers in the field of aerospace-embodied intelligence. While there are more 2D visual-language models available that are generally more mature, we focus on LLaVA[33], MiniGPT4[34], and BLIP2[35], categorizing them into $7B$ and $13B$ models based on parameter scales. Given the limited availability of open-source 3D visual-language models, we select only the 3D-LLM[9] as our research focus.
@@ VI-B Evaluation Metrics
Traditional Metrics. Common indicators include BLEU-1, BLEU-2, BLEU-3, and BLEU-4[36]. Compare the degree of overlap between the n-grams in the candidate translation and the reference translation. It is commonly used for evaluating translation quality and can be divided into multiple evaluation indicators based on n-grams.
CIDEr[37] is an evaluation metric used to assess image description tasks. Its main idea is to treat each sentence as a document, then calculate its n-gram TF-IDF vector, and use cosine similarity to measure the semantic consistency between candidate sentences and reference sentences.
SPICE [19] utilizes graph-based semantic representations to encode objects, attributes, and relationships within descriptions. Initially, it parses both the description under evaluation and the reference description into a syntactic dependency tree using a Probabilistic Context-Free Grammar (PCFG) dependency parser.
GPT4-based Metrics. GPT-4[20] has achieved significant success in aligning with human preferences. Consequently, we introduce an automated evaluation method based on GPT-4 for tasks related to aerospace embodied scene awareness, spatial reasoning, navigational exploration, and path planning. This method aims to produce evaluation results that closely resemble human assessments. By designing various prompt templates, we can effectively address different evaluation concerns.
@@
LLM-Judge-Reason&\&&Nav. The prompt language is aligned with that of llm-judge[24], enabling GPT-4 to analyze the correlation and utility between AI assistant responses and reference answers. This process aims to objectively identify and correct errors to the greatest extent possible, provide explanations, and ultimately assign scores.
@@
VII Results
VII-A Quantitative Analysis
@@
3d-llm[9] 0.0346 0.0162 0.1660 0.0283 0.0117 0.1156 0.0429 0.0197 0.1511 0.0335 0.0105 0.1378
gpt-4-vision-preview[20] 0.1200 0.0884 0.6840 0.1039 0.0924 0.6909 0.1035 0.0901 0.6511 0.1246 0.0999 0.7444
gpt-4o[20] 0.1539 0.1114 0.6800 0.1532 0.1277 0.7178 0.1237 0.1175 0.6977 0.1606 0.1225 0.7089
blip2-flan-t5-xxl[35] 0.1954 0.0860 0.4041 0.1932 0.0956 0.4333 0.2151 0.0906 0.4318 0.2593 0.1307 0.5089
blip2-opt-6.7b[35] 0.1968 0.0814 0.4201 0.2279 0.0836 0.4289 0.2140 0.0906 0.4533 0.2558 0.1102 0.4778
instructblip-flan-t5-xxl[38] 0.2118 0.0808 0.4908 0.1972 0.0852 0.4689 0.2149 0.0969 0.5067 0.2536 0.1202 0.5400
instructblip-vicuna-7b[38] 0.2239 0.0787 0.4911 0.2102 0.0835 0.5022 0.2109 0.0867 0.4889 0.2729 0.1193 0.5511
instructblip-vicuna-13b[38] 0.2185 0.0810 0.4752 0.2176 0.0852 0.4533 0.2161 0.0832 0.4556 0.2715 0.1084 0.5644
llama-adapter-v2-7B[39] 0.0843 0.0512 0.5236 0.0741 0.0546 0.5067 0.0730 0.0584 0.5378 0.0981 0.0715 0.5778
llava-v1.5-vicuna-7b[32] 0.0746 0.0469 0.5000 0.0639 0.0515 0.5364 0.0590 0.0533 0.5133 0.0790 0.0645 0.5933
llava-v1.5-vicuna-13b[32] 0.0731 0.0468 0.5314 0.0643 0.0545 0.5727 0.0604 0.0505 0.5511 0.0754 0.0673 0.6067
llava-v1.6-vicuna-7b[32] 0.0483 0.0019 0.4823 0.0423 0.0025 0.5289 0.0387 0.0036 0.5178 0.0545 0.0148 0.5778
llava-v1.6-vicuna-13b[32] 0.0484 0.0039 0.5010 0.0417 0.0024 0.5364 0.0395 0.0047 0.4738 0.0525 0.0057 0.5489
minigpt4[34] 0.0969 0.0613 0.5592 0.0824 0.0584 0.5467 0.0787 0.0605 0.4607 0.0801 0.0624 0.4489
mplug[40][40] 0.0605 0.0450 0.5626 0.0520 0.0490 0.5533 0.0522 0.0489 0.5585 0.0680 0.0582 0.5400
mplug2[41] 0.0928 0.0502 0.5276 0.0825 0.0590 0.5796 0.0675 0.0447 0.5705 0.1020 0.0714 0.5614
qwen-lv-7b[42] 0.2305 0.0946 0.4780 0.2417 0.1136 0.5133 0.2206 0.0946 0.4467 0.2682 0.1057 0.4733
@@
3d-llm[9] 0.1310 0.1008 0.3180 0.1839 0.1305 0.3133 0.0532 0.0373 0.1778 0.0792 0.009 0.2889
gpt-4-vision-preview[20] 0.0696 0.0701 0.3680 0.0830 0.1233 0.4578 0.0261 0.0154 0.3600 0.0917 0.1064 0.2822
gpt-4o[20] 0.1498 0.1710 0.493 0.1809 0.2034 0.4733 0.0558 0.0608 0.4467 0.3213 0.3750 0.4844
blip2-flan-t5-xxl[35] 0.0661 0.0863 0.3387 0.0867 0.1252 0.2756 0.0174 0.0089 0.1978 0.0868 0.0677 0.3844
blip2-opt-6.7b[35] 0.0508 0.0685 0.2023 0.0452 0.0804 0.1444 0.0533 0.0405 0.2089 0.0548 0.1619 0.2156
instructblip-flan-t5-xxl[38] 0.0966 0.1207 0.3590 0.1351 0.1725 0.2556 0.0354 0.0296 0.2133 0.1261 0.0857 0.3800
instructblip-vicuna-7b[38] 0.0254 0.0393 0.2630 0.0207 0.0493 0.2244 0.0480 0.0649 0.2978 0.0865 0.1088 0.2667
instructblip-vicuna-13b[38] 0.0158 0.0116 0.3620 0.0260 0.0278 0.2867 0.0041 0 0.1978 0.0002 0 0.2800
llama-adapter-v2-7B[39] 0.1582 0.1792 0.3430 0.1720 0.2164 0.3822 0.0721 0.0702 0.2422 0.3137 0.4432 0.5068
llava-v1.5-vicuna-7b[32] 0.1054 0.1269 0.3380 0.1190 0.1587 0.3046 0.0422 0.0412 0.4000 0.2339 0.3033 0.3667
llava-v1.5-vicuna-13b[32] 0.1235 0.1386 0.3760 0.1205 0.1837 0.3489 0.0509 0.0473 0.3911 0.2159 0.2779 0.3600
llava-v1.6-vicuna-7b[32] 0.0653 0.0887 0.3020 0.1016 0.1517 0.3444 0.0214 0.0196 0.3046 0.1123 0.1417 0.2733
llava-v1.6-vicuna-13b[32] 0.0680 0.0969 0.3490 0.0731 0.1176 0.3178 0.0250 0.0275 03378 0.106 0.164 0.3556
minigpt4[34] 0.1307 0.1714 0.3930 0.1211 0.1895 0.260 0.0288 0.0306 0.3022 0.1784 0.2266 0.3556
mplug[40] 0.1277 0.1436 0.313 0.1551 0.1932 0.3133 0.0659 0.0770 0.3711 0.2356 0.3148 0.3156
mplug2[41] 0.1375 0.1303 0.373 0.1468 0.1444 0.3288 0.0649 0.0520 0.3800 0.2668 0.238 0.4222
qwen-lv-7b[42] 0.1310 0.1590 0.305 0.1475 0.1878 0.2932 0.0873 0.0719 0.6244 0.2432 0.3324 0.3796

### v2

@@ V Dataset Suite
Figure 4: The left and right panels illustrate the construction schemes and statistics of the AerialAgent-Ego15k and CyberAgent-Ego500k datasets, respectively.
@@
V-A AerialAgent-Ego15k
Multi-Resolution UAV First-Person View City Images. The first-person view images of real cities captured by drones are derived from the UrbanBIS dataset, which is collected using aerial photogrammetry and encompasses a wide array of urban scenes. Specifically, the UrbanBIS dataset[Yang2023UrbanBISAL] comprises $0.5$ TB of aerial photographs from six actual locations: Qingdao, Wuhu, Longhua, Yuehai, Lihu, and Yingrenshi, covering a significant urban area of $10.78$ km² and including $3,370$ buildings, with a total of $113,346$ aerial photogrammetry images. We have requested images from the authors for the regions of Lihu, Longhua, Yingrenshi, and Yuehai, with resolutions of $6000\times 4000$, $8192\times 5640$, $5472\times 3648$, and $5472\times 3648$, respectively, yielding a total of $15,094$ images. From this dataset, we randomly selected $10,000$ images to serve as first-person view representations of real cities captured by drones.
Fine-grained Multi-attribute First-view Text Generation. To generate high-quality environmental descriptions, we utilize LLaVA-1.5-13B[Liu2023ImprovedBW] to produce detailed accounts of surrounding buildings, roads, trees, and other scenery from first-person perspective images captured by a drone, as illustrated in Figure 4 left (a). To standardize the format of the environmental descriptions generated by LLaVA-1.5-13B[Liu2023ImprovedBW], we employ specific prompts that emphasize the quantity, appearance, and shape of the buildings in the images, particularly focusing on the spatial relationships among the objects. This approach enhances the spatial reasoning capabilities of the drone agent. Furthermore, we specify that the sky should not be described, as this scene is relatively uniform and appears consistent from various perspectives of the drone, providing insufficient information. Consequently, the generated descriptions ensure a degree of diversity, accuracy, and detail.
Diverse Data Distribution. We perform a quantitative statistical analysis on AerialAgent-Ego15k. Figure 4 (b) and (c) illustrate the probability density functions (PDFs) of text vocabulary length and text sentence length, respectively, both exhibiting a shape akin to a normal distribution. This finding supports the rationality of the text distribution. The maximum length for image descriptions is $440$ words, with an average length of $144$ words. The maximum number of sentences in image descriptions is $42$, with an average of $11$ sentences per image. Both the number of sentences and text lengths exceed those of most existing visual-language datasets. Figure 4 (d) reveals that the dataset contains a total of $158,379$ sentences and $2,167,455$ words, of which $21,489$ are unique.
@@ V-B CyberAgent-Ego500k
First-Person Image-Text-Pose Generation. As illustrated on the right side of Figure 4 (a), the dataset construction method aligns with that of AerialAgent-15k and exhibits the following three characteristics:
@@ V-C SkyAgent-Scene3k
[…]
@@ V-D SkyAgent-Reason3k
Figure 6: The SkyAgent-Reason3k dataset contains 66 types of reasoning patterns and corresponding concrete examples, as well as some statistical results.
@@ V-G SkyAgent-Act3k
Figure 9: Examples of drone agent actions in 22 city scenarios (i.e., ShangHai and School) in SkyAgent-Act3k dataset, along with some instructions and statistics.
@@
V-H Image Enhancement
To fully consider the challenges faced by drones in real-world complex environments, this study systematically addressed two key issues encountered during flight through data augmentation. First, regarding the image tilt caused by drone body jitter, we collect more diverse data on camera tilts in the environment. Second, to address image blurring or pose drift due to jitter, we simulate the motion blur degradation process of drone images by constructing directional motion blur kernels (PSF) and applying convolutional operations. Based on the assumption of linear uniform motion, the algorithm first generates blur kernels of specified size and angle, ensuring energy conservation through normalization, and then employs two-dimensional discrete convolution to achieve image degradation. This method effectively simulates typical motion blur caused by drone body jitter or pose drift, as shown in figure 10.
Figure 10: Figure (a) and (b) show the motion blur effects in the horizontal direction and the 45-degree direction, respectively.
@@ VI-A Baselines
Baselines Selection. Due to the current scarcity of research on aerospace-embodied world models, we evaluate several mainstream and representative 3D and 2D visual-language models. This assessment aims to explore their potential and limitations concerning the proposed aerospace-embodied downstream task datasets, thereby providing a preliminary foundation for future researchers in the field of aerospace-embodied intelligence. While there are more 2D visual-language models available that are generally more mature, we focus on LLaVA[Liu2023VisualIT], MiniGPT4[Zhu2023MiniGPT4EV], and BLIP2[Li2023BLIP2BL], categorizing them into $7B$ and $13B$ models based on parameter scales. Given the limited availability of open-source 3D visual-language models, we select only the 3D-LLM[Hong20233DLLMIT] as our research focus.
@@ VI-B Evaluation Metrics
Traditional Metrics. Common indicators include BLEU-1, BLEU-2, BLEU-3, and BLEU-4[2002BLEU]. Compare the degree of overlap between the n-grams in the candidate translation and the reference translation. It is commonly used for evaluating translation quality and can be divided into multiple evaluation indicators based on n-grams.
CIDEr[Vedantam2014CIDErCI] is an evaluation metric used to assess image description tasks. Its main idea is to treat each sentence as a document, then calculate its n-gram TF-IDF vector, and use cosine similarity to measure the semantic consistency between candidate sentences and reference sentences.
SPICE [Anderson2016SPICESP] utilizes graph-based semantic representations to encode objects, attributes, and relationships within descriptions. Initially, it parses both the description under evaluation and the reference description into a syntactic dependency tree using a Probabilistic Context-Free Grammar (PCFG) dependency parser.
GPT4-based Metrics. GPT-4[Achiam2023GPT4TR] has achieved significant success in aligning with human preferences. Consequently, we introduce an automated evaluation method based on GPT-4 for tasks related to aerospace embodied scene awareness, spatial reasoning, navigational exploration, and path planning. This method aims to produce evaluation results that closely resemble human assessments. By designing various prompt templates, we can effectively address different evaluation concerns.
@@
LLM-Judge-Reason&\&Nav. The prompt language is aligned with that of llm-judge[Zheng2023JudgingLW], enabling GPT-4 to analyze the correlation and utility between AI assistant responses and reference answers. This process aims to objectively identify and correct errors to the greatest extent possible, provide explanations, and ultimately assign scores.
@@
Human Evaluation. While automated metrics provide scalable and consistent evaluation, they may not fully capture the nuances of human judgment, especially for complex cognitive tasks like spatial reasoning and path planning where logical coherence and real-world feasibility are paramount. To provide a more holistic assessment and to validate our automated findings, we complement our quantitative analysis with targeted human evaluation. Given the large scale of our benchmark, which makes comprehensive manual annotation infeasible, we conduct our evaluation on a randomly sampled subset of 100 instances for each of the four downstream tasks. We enlist experts to score the model-generated responses on a scale of 0 to 1 based on the following task-specific criteria: For Scene Awareness, evaluators assess the accuracy and completeness of object descriptions, focusing on attributes like color, shape, and quantity. For Spatial Reasoning, scoring is based on the logical correctness of the inferred relationships between objects. For Navigational Exploration, judgment considers both the accuracy of following the navigation command and the correctness of the answer provided. For Task Planning, scores are assigned based on the plan’s feasibility, the correctness of the action sequence, and the accuracy of landmark identification.
VII SkyAgentX: unifying perception, reasoning, planning, and navigating
Figure 12: Overview of the UAV-agent embodied large model, SkyAgentX, which integrates “Perception-Reasoning-Navigating-Planning” into an unified framework with aerospace embodied chain-of-thought and multitask curriculum learning.
VII-A Aerospace Embodied Chain-of-thought
Chain-of-thought is a mechanism that mimics the human step-by-step reasoning process. By breaking down complex problems into multiple intermediate steps, it guides the model to generate coherent and reasonable answers. Chain-of-thought explicitly displays the model’s reasoning path, making the decision-making process transparent and facilitating the analysis of error sources. For example, in drone mission planning, the model can first identify environmental features, then evaluate feasible paths, and finally generate action sequences, with each step being traceable, thereby enhancing the model’s interpretability. By solving sub-problems step by step, the model can more efficiently handle multimodal, long-sequence, or highly abstract tasks.
In the AeroVerse benchmark, we automatically expand data through large models (such as GPT-4), generate diverse reasoning paths through prompt engineering, and construct aerospace embodied Chain-of-thought training data.
VII-B Aerospace Embodied Multitask Curriculum Learning
In the AeroVerse benchmark, the design of the five downstream tasks (aerospace scene perception, spatial reasoning, navigational exploration, task planning, and motion decision-making) follows a progressive relationship from simple to complex and from perception to decision-making. This structure naturally lends itself to the training strategy of Curriculum Learning, which involves phased, incremental task training to gradually enhance the comprehensive capabilities of the drone agent.
As the starting point of Curriculum Learning, Aerospace Embodied Scene Awareness requires the drone to describe its surroundings (e.g., buildings, roads, trees, etc.) from a first-person perspective. The goal of this stage is to equip the model with basic environmental understanding capabilities. This phase resembles the “observation and description” stage in human learning, providing the foundation of environmental cognition for subsequent tasks. Through extensive scene description training, the model establishes preliminary representational abilities for three-dimensional space.
After mastering scene perception, the model must further understand the spatial relationships between objects. The Aerospace Embodied Spatial Reasoning task requires the drone to answer questions about the environment, upgrading from static descriptions to dynamic reasoning. The model must perform logical judgments by combining three-dimensional scene features (e.g., depth information, camera pose). This stage reinforces the model’s spatial modeling and causal reasoning capabilities through complex questions in the SkyAgent-Reason3k dataset, laying the groundwork for subsequent navigation tasks. The Aerospace Embodied Navigational Exploration task requires the drone to explore the environment based on long-distance navigation instructions (e.g., “Fly forward 200 meters and then turn left”) and answer attribute-related questions about objects encountered along the way. The navigation task introduces a temporal dimension and action sequences, requiring the model to translate the perception and reasoning abilities learned in the previous stages into concrete actions. Through training on the SkyAgent-Nav3k dataset, the model learns to achieve a preliminary “perception-reasoning-action” loop in complex urban environments.
The Aerospace Embodied Task Planning task is positioned as the advanced planning stage in the curriculum. This task requires the drone to generate detailed path plans (e.g., “First go straight to the red building, then turn right and bypass the tall building”) based on the starting and ending points. This stage is an extension of the navigation task but places greater emphasis on global planning capabilities.
VII-C SkyAgent & SkyAgentX
As illustrated in Figure 12, based on a pre-trained general vision-language model Internvl-2.5-8B[chen2024internvl], we introduce the aerospace embodied chain-of-thought mechanism and fine-tune it using specialized instruction datasets (i.e., SkyAgent-Scene3k, SkyAgent-Reason3k, SkyAgent-Nav3k, and SkyAgent-Plan3k) for four downstream tasks, to develop the SkyAgent model. On this foundation, we further incorporate the aerospace embodied multitask curriculum learning strategy, progressively training the model through multi-task joint training in the sequential order of scene perception, spatial reasoning, navigation exploration, and task planning. This process ultimately led to the construction of the UAV-agent embodied large model, SkyAgentX, which integrates “perception-reasoning-navigating-planning” into an unified framework.
VIII Results
VIII-A Quantitative Analysis
@@
3d-llm[Hong20233DLLMIT] 0.0346 0.0162 0.1660 0.0283 0.0117 0.1156 0.0429 0.0197 0.1511 0.0335 0.0105 0.1378
GPT-4-vision-preview[Achiam2023GPT4TR] 0.1200 0.0884 0.6840 0.1039 0.0924 0.6909 0.1035 0.0901 0.6511 0.1246 0.0999 0.7444
GPT-4o[Achiam2023GPT4TR] 0.1539 0.1114 0.6800 0.1532 0.1277 0.7178 0.1237 0.1175 0.6977 0.1606 0.1225 0.7089
blip2-flan-t5-xxl[Li2023BLIP2BL] 0.1954 0.0860 0.4041 0.1932 0.0956 0.4333 0.2151 0.0906 0.4318 0.2593 0.1307 0.5089
blip2-opt-6.7b[Li2023BLIP2BL] 0.1968 0.0814 0.4201 0.2279 0.0836 0.4289 0.2140 0.0906 0.4533 0.2558 0.1102 0.4778
instructblip-flan-t5-xxl[Dai2023InstructBLIPTG] 0.2118 0.0808 0.4908 0.1972 0.0852 0.4689 0.2149 0.0969 0.5067 0.2536 0.1202 0.5400
instructblip-vicuna-7b[Dai2023InstructBLIPTG] 0.2239 0.0787 0.4911 0.2102 0.0835 0.5022 0.2109 0.0867 0.4889 0.2729 0.1193 0.5511
instructblip-vicuna-13b[Dai2023InstructBLIPTG] 0.2185 0.0810 0.4752 0.2176 0.0852 0.4533 0.2161 0.0832 0.4556 0.2715 0.1084 0.5644
llama-adapter-v2-7B[Gao2023LLaMAAdapterVP] 0.0843 0.0512 0.5236 0.0741 0.0546 0.5067 0.0730 0.0584 0.5378 0.0981 0.0715 0.5778
llava-v1.5-vicuna-7b[Liu2023ImprovedBW] 0.0746 0.0469 0.5000 0.0639 0.0515 0.5364 0.0590 0.0533 0.5133 0.0790 0.0645 0.5933
llava-v1.5-vicuna-13b[Liu2023ImprovedBW] 0.0731 0.0468 0.5314 0.0643 0.0545 0.5727 0.0604 0.0505 0.5511 0.0754 0.0673 0.6067
llava-v1.6-vicuna-7b[Liu2023ImprovedBW] 0.0483 0.0019 0.4823 0.0423 0.0025 0.5289 0.0387 0.0036 0.5178 0.0545 0.0148 0.5778
llava-v1.6-vicuna-13b[Liu2023ImprovedBW] 0.0484 0.0039 0.5010 0.0417 0.0024 0.5364 0.0395 0.0047 0.4738 0.0525 0.0057 0.5489
miniGPT4[Zhu2023MiniGPT4EV] 0.0969 0.0613 0.5592 0.0824 0.0584 0.5467 0.0787 0.0605 0.4607 0.0801 0.0624 0.4489
mplug[Ye2023mPLUGOwlME][Ye2023mPLUGOwlME] 0.0605 0.0450 0.5626 0.0520 0.0490 0.5533 0.0522 0.0489 0.5585 0.0680 0.0582 0.5400
mplug2[Ye2023mPLUGOwl2RM] 0.0928 0.0502 0.5276 0.0825 0.0590 0.5796 0.0675 0.0447 0.5705 0.1020 0.0714 0.5614
qwen-lv-7b[Bai2023QwenVLAV] 0.2305 0.0946 0.4780 0.2417 0.1136 0.5133 0.2206 0.0946 0.4467 0.2682 0.1057 0.4733
SkyAgent (ours) 0.4302 0.3083 0.5166 0.5349 0.4068 0.4831 0.5085 0.3349 0.4732 0.4278 0.2557 0.5295
@@
3d-llm[Hong20233DLLMIT] 0.1310 0.1008 0.3180 0.1839 0.1305 0.3133 0.0532 0.0373 0.1778 0.0792 0.009 0.2889
GPT-4-vision-preview[Achiam2023GPT4TR] 0.0696 0.0701 0.3680 0.0830 0.1233 0.4578 0.0261 0.0154 0.3600 0.0917 0.1064 0.2822
GPT-4o[Achiam2023GPT4TR] 0.1498 0.1710 0.493 0.1809 0.2034 0.4733 0.0558 0.0608 0.4467 0.3213 0.3750 0.4844
blip2-flan-t5-xxl[Li2023BLIP2BL] 0.0661 0.0863 0.3387 0.0867 0.1252 0.2756 0.0174 0.0089 0.1978 0.0868 0.0677 0.3844
blip2-opt-6.7b[Li2023BLIP2BL] 0.0508 0.0685 0.2023 0.0452 0.0804 0.1444 0.0533 0.0405 0.2089 0.0548 0.1619 0.2156
instructblip-flan-t5-xxl[Dai2023InstructBLIPTG] 0.0966 0.1207 0.3590 0.1351 0.1725 0.2556 0.0354 0.0296 0.2133 0.1261 0.0857 0.3800
instructblip-vicuna-7b[Dai2023InstructBLIPTG] 0.0254 0.0393 0.2630 0.0207 0.0493 0.2244 0.0480 0.0649 0.2978 0.0865 0.1088 0.2667
instructblip-vicuna-13b[Dai2023InstructBLIPTG] 0.0158 0.0116 0.3620 0.0260 0.0278 0.2867 0.0041 0 0.1978 0.0002 0 0.2800
llama-adapter-v2-7B[Gao2023LLaMAAdapterVP] 0.1582 0.1792 0.3430 0.1720 0.2164 0.3822 0.0721 0.0702 0.2422 0.3137 0.4432 0.5068
llava-v1.5-vicuna-7b[Liu2023ImprovedBW] 0.1054 0.1269 0.3380 0.1190 0.1587 0.3046 0.0422 0.0412 0.4000 0.2339 0.3033 0.3667
llava-v1.5-vicuna-13b[Liu2023ImprovedBW] 0.1235 0.1386 0.3760 0.1205 0.1837 0.3489 0.0509 0.0473 0.3911 0.2159 0.2779 0.3600
llava-v1.6-vicuna-7b[Liu2023ImprovedBW] 0.0653 0.0887 0.3020 0.1016 0.1517 0.3444 0.0214 0.0196 0.3046 0.1123 0.1417 0.2733
llava-v1.6-vicuna-13b[Liu2023ImprovedBW] 0.0680 0.0969 0.3490 0.0731 0.1176 0.3178 0.0250 0.0275 03378 0.106 0.164 0.3556
miniGPT4[Zhu2023MiniGPT4EV] 0.1307 0.1714 0.3930 0.1211 0.1895 0.260 0.0288 0.0306 0.3022 0.1784 0.2266 0.3556
mplug[Ye2023mPLUGOwlME] 0.1277 0.1436 0.313 0.1551 0.1932 0.3133 0.0659 0.0770 0.3711 0.2356 0.3148 0.3156
mplug2[Ye2023mPLUGOwl2RM] 0.1375 0.1303 0.373 0.1468 0.1444 0.3288 0.0649 0.0520 0.3800 0.2668 0.238 0.4222
qwen-lv-7b[Bai2023QwenVLAV] 0.1310 0.1590 0.305 0.1475 0.1878 0.2932 0.0873 0.0719 0.6244 0.2432 0.3324 0.3796
SkyAgent (ours) 0.4598 0.3846 0.4310 0.3552 0.3255 0.2729 0.3323 0.1648 0.5792 0.3570 0.3739 0.3189
SkyAgentX (ours) 0.4660 0.4185 0.5309 0.3601 0.3195 0.3295 0.3469 0.1777 0.6043 0.4399 0.4439 0.4197